If one solution to the equation x^2 – 3x + k = 0 is 4, then what is the other solution?
a) -4 b) -1 c) 1 d) 4 e) cannot be determined from the information given

Answer by rapaljer(4671)

Substitute x=4 and solve for k:
4^2-3*4+k=0
16-12+k=0
4+k=0
k=-4

Dr. Robert J. Rapalje, Retired